Show your work and include units with your answers for full credit.

The “A” string on a guitar is 65 cm long and has a fourth harmonic whose frequency is f4 = 880 Hz.

(a) What is the fundamental frequency (f1) of the string?

(b) What is the speed of waves on the string?

(c) What is the wavelength of the fundamental waves you found in part (a)?
Answer #1

a) fundamental frequency is f1

fourth harmonic frequency is f4 = 4*f1 = 880 Hz

f1 = 880/4 = 220 Hz

b) f1 = v/(2*l)


speed is v = f1*2*l = 220*2*0.65 = 286 m/s

c) v = lamda*f


286 = lamda*220

lamda = 286/220 = 1.3 m